OpenStudy (anonymous):

Using natural log, solve for x. 2^x-4 = 5.

OpenStudy (anonymous):

Can someone please show me the steps

OpenStudy (aum):

Is it \(\large 2^{x-4} = 5\) ?

OpenStudy (anonymous):

yes

OpenStudy (aum):

\[ \large 2^{x-4} = 5 \\ \text{ } \\ \text{Take the natural logarithm on both sides: } \\ \text{ } \\ \ln(2^{x-4}) = \ln(5) \\ \large (x-4)\ln(2) = \ln(5) \\ x - 4 = \frac{\ln(5)}{\ln(2)} \\ x = 4 + \frac{\ln(5)}{\ln(2)} \]

OpenStudy (anonymous):

thank you!